# Phase transitions and remnants of fractionalization at finite temperature in the triangular lattice quantum loop model

### v3

@@ Abstract
Figure 1: Finite-temperature phase diagram of the QLM on the triangular lattice which exhibits different phases: Lattice Nematic (LN, green region), Vison Plaquette (VP, blue region), and disordered (red) phases. The solid gray horizontal line marks the $\mathbb{Z}_{2}$ spin liquid phase only existing in the ground-state. The green $(V\simeq 0.05)$ and blue $(V\simeq 0.6)$ points correspond to the LN-VP and VP-QSL quantum phase transitions. The red star indicates a possible multicritical point at $(V,T)=(0.07(2),0.22(2))$. The two dashed gray lines correspond to $T=0.23$ (upper) and $T=0.1$ (lower) used in Fig. 4. The inset zooms in at low-$T$ near the VP-QSL Cubic* QCP, highlighting that the finite-$T$ phase boundary scales as $T_{c}(V)\sim|V-V_{c}|^{\nu_{C^{*}}}$, with $\nu_{C^{*}}\simeq 0.7$ the correlation length exponent for the 3d Cubic* universality class.
Simulations and Order Parameters.— We study Eq. 1 using QMC simulations for system sizes $N=L^{2}$ with $L=8,12,16,20,24,28,32$. On top of lattice symmetries, the model Eq. (1) exhibits topological sectors labeled by the parity of the winding number of loops. We restrict our QMC simulations in the $(0,0)$ topological sector to decrease statistical errors, but we checked using a fully-ergodic extension of the sweeping cluster algorithm Dabholkar et al. (2022) as well as exact diagonalization data (see Supplemental Material (SM) sup ) that this does not affect our conclusions as the finite-$T$ properties discussed below are dominated by the $(0,0)$ sector.
@@
[…]
Figure 2: Left panels: data collapse for (a) the nematic order parameter $\langle D\rangle$, (b) its Binder ratio $B_{D}$, (c) the vison parameter $\phi$ and (d) its Binder ratio $B_{|\phi|}$ for system sizes from $L=8$ to $L=24$. Simulations are performed at $V=-0.5$, where the ground state is deep within the LN phase. All data collapses for $\langle D\rangle$, $B_{D}$, $\phi$ and $B_{|\phi|}$ use the critical exponents $\beta_{3P}=\frac{1}{9}$ and $\nu_{3P}=\frac{5}{6}$ of the 3-state Potts model. Right panels: (e) Binder ratio $B_{|\phi|}$ of the vison order parameter and (f) Correlation ratio $R_{C{\cal T}}$ of the $t$-terms for system sizes $L=4$ to $L=32$ for $V=0.3$ (where the ground state is located within the VP phase). The inset illustrates the crossing between system sizes $L$ and $2L$ for both $B_{|\phi|}$ and $R_{C{\cal T}}$. Lines are fits to $T_{c}(L)=T_{c}(\infty)+aL^{-(\omega+1/\nu)}$, where $\omega$ is the correction exponent and $\nu$ the correlation length critical exponent. For $V=0.3$, we obtain $T_{c}(\infty)=0.14(1)$ and $\omega+1/\nu\sim 1.4$. Similar analysis for other $V$ values are presented in SM sup .
@@
LN-Disorder 3-State Potts Transition.— Starting on the left part of the phase diagram, we obtain a very good finite-size collapse of the nematic order parameter $\langle D\rangle$ and its Binder ratio $B_{D}=\langle D^{4}\rangle/\langle D^{2}\rangle^{2}$ in Fig. 2 (a) and (b) at $V=-0.5$ using the critical exponents of the 3-state Potts universality class $\beta_{3P}=\frac{1}{9}$ and $\nu_{3P}=\frac{5}{6}$ Baxter (2007); Wu (1982); Alexander (1975); Baxter (2007); Wu (1982), with the interpretation that the three states in the Potts model correspond to the possible orientations of loops in the LN phase. The empty triangles in Fig. 1 denote the LN-Disorder transitions at other values of $V$ as determined from a similar analysis. In addition to the nematic order parameter, we also monitor the critical behavior of the vison modes $\phi$ for the same LN-Disorder transition. The data collapse of $|\phi|$ and its Binder ratio $B_{|\phi|}={\langle|\bm{\phi}|^{4}\rangle}/{\langle|\bm{\phi}|^{2}\rangle^{2}}$ with $|\phi|=\sqrt{\phi_{1}^{2}+\phi_{2}^{2}+\phi_{3}^{2}}$ at $V=-0.5$ are shown in Fig. 2 (c) and (d) with the same exponents as for the nematic order parameter. This can be intuitively understood as the existence of a symmetry-breaking pattern in the loop configurations (leading to a finite expectation value for $\langle D\rangle$) induces in a deterministic fashion a similar symmetry-breaking for the $\phi$ field (see SM sup for a schematic presentation of $\phi$ and loop-segment fields) with both 3-fold rotation and translation being broken. Therefore $D$ or $|\phi|$ are the same order parameter for the LN-Disorder transition, and there is no remnant of fractionalization at finite-$T$ in this part of the phase diagram.
@@
VP-Disorder Transition and “remnants of fractionalization" at finite-$T$.— The transition out of the intrinsically quantum VP phase is more subtle as it cannot be easily captured by a classical constrained model with the same symmetries (as e.g. a classical dimer model captures the finite-$T$ columnar transition in the square lattice QDM Dabholkar et al. (2022)). We first discuss theoretical expectations for the VP-Disorder transition before comparing to numerical simulations. Near the QCP at $V_{c}\simeq 0.6$ of Eq. (2), the critical temperature should follow a power law behavior $T_{c}\sim|V-V_{c}|^{\nu_{C^{*}}}$, with $\nu_{C^{*}}$ the critical exponent of the Cubic* QCP (for more details see SM sup ). The finite-temperature VP-Disorder transition has two possible interpretations. Putting Eq. (2) at finite $T$, we expect a scalar theory respecting the Cubic symmetry in two spatial dimensions. One candidate theory is three decoupled copies of Ising models since the Cubic conformal field theory (CFT) merge with the decoupled Ising models at $D=2$. Away from the quantum critical point, however, another candidate is the 4-state Potts universality class. In this case, the gauge invariant order parameter for the VP phase is given by the kinetic $t_{i}$-terms, which becomes the order parameter of the 4-state Potts model. The details of the above discussion are explained in SM sup . The crucial difference between these two interpretations are whether the kinetic $t_{i}$-terms and the vison field $\phi$ have different scaling dimensions at criticality.
@@
In this Supplemental Material, we first present the review of the mapping between the dimer and vison configurations and the ground-state phase diagram for the QLM on the triangular lattice in Sec. I. Then, the results of exact diagonalization for QLM on the triangular lattice are shown in Sec. II. In Sec. III, we discuss finite-size scaling and data collapse through a Bayesian scaling analysis of the order parameters in the VP-Disorder phase transition. Thereafter, the determination of the LN-VP transition boundary and the estimation of the multicritical point are performed through an analysis of the histograms of $\phi$ in Sec. IV. Finally, Sec. V offers a field-theoretical discussion on the finite-temperature VP-Disorder transitions.
@@ III Finite-size scaling analysis and data collapse for the VP-Disorder phase transition
The fit results for three values of $V$ within the VP phase ($V=0.2$, $V=0.3$ and $V=0.4$) are summarized in Table. S1. When comparing these results with the critical exponents of the two-dimensional Ising model, 4-state Potts model and related Ashkin-Teller (AT) model listed in Table. S2, we first observe that our values for $\beta^{\phi}$ and $\nu$ lie within the range of critical exponents characteristic of the AT model. In the AT model, while the critical exponents $\beta$ and $\nu$ vary along the critical line, their ratio remains fixed. However, our data yield a ratio of $\beta^{\phi}$ and $\nu$ of approximately $1/4$, which differs from the $1/8$ ratio of the AT model. Furthermore, our critical exponents derived from the data collapse differ substantially from those of the 4-state Potts model or Ising model. On the other hand, we also observe that the values of $2\beta^{C{\cal T}}/\nu$ (obtained from the data collapse for $C{\cal T}$ and which corresponding to the scaling dimension of the kinetic ${\cal T}$ term) are quite large $\sim 1.4$ for $V=0.2,0.3$ and $0.4$, and close to the scaling dimension of the rank-2 symmetric traceless tensor for the 3d Cubic* and O(3)* transition. This seems to suggest that the observed finite-temperature phase boundary of VP-Disorder transition, still experiences the influence of the ground state VP-QSL 3d Cubic* QCP. It is important to note that due to the limitations in system size in our simulations, the quality of the extrapolation and data collapse are not optimal.
Table S1: Fitting results for order parameters in the VP phase
$V=0.2$ Extrapolation Data collapse for $|\phi|$ Data collapse for $B_{|\phi|}$ Data collapse for $C{\cal T}$ Data collapse for $R_{C{\cal T}}$
$T_{c}$ 0.176(7) 0.199(2) 0.206(2) 0.204(1) 0.219(3)
@@
Ising $\frac{1}{8}$ $1$ $\frac{1}{8}$
@@ IV Order parameter and histogram in the LN-VP phase transition
V Theoretical expectations for the finite-temperature VP-Disorder transition
@@
[…]

### v4

@@ Abstract
Figure 1: Finite-temperature phase diagram of the QLM on the triangular lattice which exhibits different phases: Lattice Nematic (LN, green region), Vison Plaquette (VP, blue region), and disordered (red) phases. A comprehensive summary of the data can be found in Table. S1 and S3 of the Supplemental Material (SM) sup . The solid gray horizontal line marks the $\mathbb{Z}_{2}$ spin liquid phase only existing in the ground-state. The green $(V\simeq 0.05)$ and blue $(V\simeq 0.6)$ points correspond to the LN-VP and VP-QSL quantum phase transitions. The red star indicates a possible multicritical point at $(V,T)=(0.07(2),0.22(2))$. The two dashed gray lines correspond to $T=0.23$ (upper) and $T=0.1$ (lower) used in Fig. 4. The inset zooms in at low-$T$ near the VP-QSL Cubic* QCP, highlighting that the finite-$T$ phase boundary scales as $T_{c}(V)\sim|V-V_{c}|^{\nu_{C^{*}}}$, with $\nu_{C^{*}}\simeq 0.7$ the correlation length exponent for the 3d Cubic* universality class.
Simulations and Order Parameters.— We study Eq. 1 using QMC simulations for system sizes $N=L^{2}$ with $L=8,12,16,20,24,28,32$. On top of lattice symmetries, the model Eq. (1) exhibits topological sectors labeled by the parity of the winding number of loops. We restrict our QMC simulations in the $(0,0)$ topological sector to decrease statistical errors, but we checked using a fully-ergodic extension of the sweeping cluster algorithm Dabholkar et al. (2022) as well as exact diagonalization data (see SM sup ) that this does not affect our conclusions as the finite-$T$ properties discussed below are dominated by the $(0,0)$ sector.
@@
Figure 2: Left panels: data collapse for (a) the nematic order parameter $\langle D\rangle$, (b) its Binder ratio $B_{D}$, (c) the vison parameter $\phi$ and (d) its Binder ratio $B_{|\phi|}$ for system sizes from $L=8$ to $L=24$. Simulations are performed at V=0.5V=-0.5, where the ground state is deep within the LN phase. All data collapses for $\langle D\rangle$, $B_{D}$, $\phi$ and $B_{|\phi|}$ use the critical exponents $\beta_{3P}=\frac{1}{9}$ and $\nu_{3P}=\frac{5}{6}$ of the 3-state Potts model. Right panels: (e) Binder ratio $B_{|\phi|}$ of the vison order parameter and (f) Correlation ratio $R_{C{\cal T}}$ of the $t$-terms for system sizes $L=4$ to $L=32$ for $V=0.3$ (where the ground state is located within the VP phase). The inset illustrates the crossing between system sizes $L$ and $2L$ for both $B_{|\phi|}$ and $R_{C{\cal T}}$. Lines are fits to $T_{c}(L)=T_{c}(\infty)+aL^{-(\omega+1/\nu)}$, where $\omega$ is the correction exponent and $\nu$ the correlation length critical exponent. For $V=0.3$, we obtain $T_{c}(\infty)=0.14(1)$ and $\omega+1/\nu\sim 1.4$. Similar analysis for other $V$ values are presented in SM sup .
@@
LN-Disorder 3-State Potts Transition.— Starting on the left part of the phase diagram, we obtain a very good finite-size collapse of the nematic order parameter $\langle D\rangle$ and its Binder ratio $B_{D}=\langle D^{4}\rangle/\langle D^{2}\rangle^{2}$ in Fig. 2 (a) and (b) at V=0.5V=-0.5 using the critical exponents of the 3-state Potts universality class $\beta_{3P}=\frac{1}{9}$ and $\nu_{3P}=\frac{5}{6}$ Baxter (2007); Wu (1982); Alexander (1975); Baxter (2007); Wu (1982), with the interpretation that the three states in the Potts model correspond to the possible orientations of loops in the LN phase. The empty triangles in Fig. 1 denote the LN-Disorder transitions at other values of $V$ as determined from a similar analysis. In addition to the nematic order parameter, we also monitor the critical behavior of the vison modes $\phi$ for the same LN-Disorder transition. The data collapse of $|\phi|$ and its Binder ratio $B_{|\phi|}={\langle|\bm{\phi}|^{4}\rangle}/{\langle|\bm{\phi}|^{2}\rangle^{2}}$ with $|\phi|=\sqrt{\phi_{1}^{2}+\phi_{2}^{2}+\phi_{3}^{2}}$ at V=0.5V=-0.5 are shown in Fig. 2 (c) and (d) with the same exponents as for the nematic order parameter. This can be intuitively understood as the existence of a symmetry-breaking pattern in the loop configurations (leading to a finite expectation value for $\langle D\rangle$) induces in a deterministic fashion a similar symmetry-breaking for the $\phi$ field (see SM sup for a schematic presentation of $\phi$ and loop-segment fields) with both 3-fold rotation and translation being broken. Therefore $D$ or $|\phi|$ are the same order parameter for the LN-Disorder transition, and there is no remnant of fractionalization at finite-$T$ in this part of the phase diagram.
@@
VP-Disorder Transition and “remnants of fractionalization" at finite-TT.— The transition out of the intrinsically quantum VP phase is more subtle as it cannot be easily captured by a classical constrained model with the same symmetries (as e.g. a classical dimer model captures the finite-$T$ columnar transition in the square lattice QDM Dabholkar et al. (2022)). We first discuss theoretical expectations for the VP-Disorder transition before comparing to numerical simulations. Near the QCP at $V_{c}\simeq 0.6$ of Eq. (2), the critical temperature should follow a power law behavior $T_{c}\sim|V-V_{c}|^{\nu_{C^{*}}}$, with $\nu_{C^{*}}$ the critical exponent of the Cubic* QCP (for more details see SM sup ). The finite-temperature VP-Disorder transition has two possible interpretations. Putting Eq. (2) at finite $T$, we expect a scalar theory respecting the Cubic symmetry in two spatial dimensions. One candidate theory is three decoupled copies of Ising models since the Cubic conformal field theory (CFT) merge with the decoupled Ising models at $D=2$. Away from the quantum critical point, however, another candidate is the 4-state Potts universality class. In this case, the gauge invariant order parameter for the VP phase is given by the kinetic $t_{i}$-terms, which becomes the order parameter of the 4-state Potts model. The details of the above discussion are explained in SM sup . The crucial difference between these two interpretations are whether the kinetic $t_{i}$-terms and the vison field $\phi$ have different scaling dimensions at criticality.
@@
In this Supplemental Material, we first present the review of the mapping between the dimer and vison configurations and the ground-state phase diagram for the QLM on the triangular lattice in Sec. I. Then, the results of exact diagonalization for QLM on the triangular lattice are shown in Sec. II. In Sec. III, we discuss finite-size scaling and data collapse through a Bayesian scaling analysis of the order parameters in the VP-Disorder phase transition. Thereafter, the determination of the LN-VP transition boundary and the estimation of the multicritical point are performed through an analysis of the histograms of $\phi$ in Sec. IV. Supplemental QMC data for the LN-disordered phase transition is presented in Sec. V. Finally, Sec. VI offers a field-theoretical discussion on the finite-temperature VP-Disorder transitions.
@@ III Finite-size scaling analysis and data collapse for the VP-Disorder phase transition
The fit results for different values of $V$ within the VP phase are summarized in Table. S1. We note that for values of $V$ close to the ground-state phase transition points, scaling becomes increasingly challenging. For instance, the data collapse for $C{\cal T}$ at $V=0.1$ fails unless we fix $T_{c}=0.2$, as indicated in Table. S1. similarly, the data collapses for $C{\cal T}$ and $R_{C{\cal T}}$ also fail at $V=0.55$. For each $V$, we computed a weighted average of $T_{c}$ from the data collapse results of different physical observables, using the Log-likelihood from each data collapse result as the weight, as indicated in Table. S1. The weighted average is calculated using $\sum(w_{i}x_{i})/\sum w_{i}$, with the errorbar given by $\sqrt{\sum(w_{i}^{2}\sigma_{i}^{2})/(\sum w_{i})^{2}}$. The final estimated $T_{c}$ in the last column of Table. S1 is chosen to balance all estimates, with an errorbar encompassing all uncertainties. The larger error bar in Fig. 1 of the main text reflects the combined uncertainty from all methods, thereby reducing bias toward any single scaling.
When comparing these results with the critical exponents of the two-dimensional Ising model, 4-state Potts model and related Ashkin-Teller (AT) model listed in Table. S2, we first observe that our values for $\beta^{\phi}$ and $\nu$ lie within the range of critical exponents characteristic of the AT model. In the AT model, while the critical exponents $\beta$ and $\nu$ vary along the critical line, their ratio remains fixed. However, our data yield a ratio of $\beta^{\phi}$ and $\nu$ of approximately $1/4$, which differs from the $1/8$ ratio of the AT model. Furthermore, our critical exponents derived from the data collapse differ substantially from those of the 4-state Potts model or Ising model. On the other hand, we also observe that the values of $2\beta^{C{\cal T}}/\nu$ (obtained from the data collapse for $C{\cal T}$ and which corresponding to the scaling dimension of the kinetic $\cal T$} term) are quite large $\sim 1.4$ for $V=0.2,0.3$ and $0.4$, and close to the scaling dimension of the rank-2 symmetric traceless tensor for the 3d Cubic* and O(3)* transition. This seems to suggest that the observed finite-temperature phase boundary of VP-Disorder transition, still experiences the influence of the ground state VP-QSL 3d Cubic* QCP. It is important to note that due to the limitations in system size in our simulations, the quality of the extrapolation and data collapse are not optimal.
Table S1: Fitting results for order parameters in the VP phase. The merged column shows the data collapse results for different observables.
$V=0.1$ Extrapolation $|\phi|$ $B_{|\phi|}$ $C{\cal T}$ $R_{C{\cal T}}$ Weighted average Estimated $T_{c}$
$T_{c}$ 0.190(8) 0.226(2) 0.247(3) Fixed 0.2 0.254(2) 0.238(2) 0.21(3)
$\beta^{\phi}$ 0.212(6)
$\beta^{C{\cal T}}$ 0.59(2)
$\nu$ 0.92(2) 0.80(6) 0.97(3) 0.92(6) 0.89(4)
$(\omega+1/\nu)_{B_{|\phi|}}$ 1.1(2)
$(\omega+1/\nu)_{R_{C{\cal T}}}$ 1.1(1)
Log-likelihood 250.8 137.4 163.8 108.2
$V=0.2$
$T_{c}$ 0.176(7) 0.199(2) 0.206(2) 0.204(1) 0.219(3) 0.206(2) 0.20(2)
@@
Log-likelihood 208 173.6 217.2 120.2
$V=0.5$
$T_{c}$ 0.093(6) 0.098(3) 0.107(2) 0.097(3) 0.111(2) 0.102(3) 0.09(2)
$\beta^{\phi}$ 0.23(2)
$\beta^{C{\cal T}}$ 0.50(2)
$\nu$ 0.73(4) 0.77(4) 0.64(3) 0.66(4) 0.70(4)
$(\omega+1/\nu)_{B_{|\phi|}}$ 1.3(3)
$(\omega+1/\nu)_{R_{C{\cal T}}}$ 1.9(2)
Log-likelihood 144.9 121.3 151.1 102.8
$V=0.55$
$T_{c}$ 0.060(9) 0.048(5) 0.087(2) 0.067(4) 0.06(2)
$\beta^{\phi}$ 0.36(2)
$\beta^{C{\cal T}}$
$\nu$ 1.01(5) 0.63(3) 0.83(4)
$(\omega+1/\nu)_{B_{|\phi|}}$
$(\omega+1/\nu)_{R_{C{\cal T}}}$ 2.0(3)
Log-likelihood 146.7 132.7
@@
Ising $\frac{1}{8}$ 11 $\frac{1}{8}$
@@ IV Order parameter and histogram in the LN-VP phase transition
V Numerical results for the LN-disordered phase transition
In this section, we present the raw QMC data for the nematic order parameter and its Binder ratio for the transition between the LN and disordered phases in Fig. S7. This data serves as supplemental information for the left panel of Fig. 2 in the main text. We conduct similar simulations and analyses for other values of $V$ and summarize the transition temperatures in Table. S3, as illustrated in the phase diagram in Fig. 1 in the main text.
Figure S7: QMC data for (a) the nematic order parameter $\langle D\rangle$ and (b) its Binder ratio as a function of temperature at V=0.5V=-0.5. The crossing points for various system sizes indicates the transition temperature $T_{c}=0.83(2)$. The errorbars are determined based on the temperature intervals in our simulations. The data collapse using the standard values of the critical exponents from the 3-state Potts universality class is shown in the left panel of Fig. 2 in the main text.
Table S3: Transition temperatures for various $V$ values in the LN-disordered phase transition.
$V$ -1 -0.5 -0.2 -0.1 0
$T_{c}$ 1.57(2) 0.83(2) 0.46(2) 0.35(2) 0.29(2)
VI Theoretical expectations for the finite-temperature VP-Disorder transition
@@
$S=\int d\tau d^{d}x\,\mathcal{L},\quad{\rm with}\quad\mathcal{L}=\frac{1}{2}\partial_{\mu}\phi\,\partial^{\mu}\phi+\frac{1}{2}\phi^{2}+\frac{1}{4!}\lambda\phi^{4}$ (S6)